# A Fast Relax-and-Round Approach to Unit Commitment with Sub Hourly Mechanical and Ramp Constraints

### v1

@@ I Introduction
Studies directed at improving computational efficiency use comparably small simulations and one hour resolution. For example, [16] uses a bundle method on a 130-generator system, [37] uses temporal bundles and 289 generators, and [30] demonstrates security constrained unit commitment on 69 generators. There are several reasons for the use of small system sizes in unit commitment problems. First, system scaling happens along both spatial and temporal granularity. Running 1000 unit commitment scenarios for a 150-unit system remains substantially faster than running one scenario for a 1000-unit system. Similarly, running 72 time decoupled unit commitment scenarios is faster than running one time coupled scenario for a three days outlook. Second, existing unit commitment algorithms rely on mixed integer programs (MIPs) [2]. In the worst case, solving MIPs is NP-hard and requires time exponential in the problem size, to enumerate all possibilities. For more information about the mathematical reasons for the poor scaling behavior of MIPs, see optimization-specific reviews such as [17, 20, 4, 28].
@@
Our approach only accounts for the costs of generation in the current period, but includes intertemporal runtime constraints and, later, ramping constraints. From that perspective, the algorithm is partially temporally decoupled and cannot be directly compared to fully coupled UC algorithms. However, it still presents an advancement over earlier methods because (i) it scales sub-quadratically rather than exponentially and (ii) does not linearize the problem, making solutions more accurate. Each extended formulation is evaluated from two viewpoints: algorithmic runtime and the objective function. For the relaxed unit commitment formulation, we further analyze the stochastic characteristics of the objective function.
The study is organized as follows. Section II summarizes the original model. Section III introduces intraday unit constraints. Section IV extends the model to include ramping constraints. Section V briefly discusses results with a different ramping formulation. Section VI summarizes the results and suggests further research directions.
@@ II-A Mathematical Formulation
  • $\sigma_{D}$: standard deviation of demand forecast
@@
The constraint Eq. 1c ensures that we supply the projected demand. The constraint Eq. 1d ensures that if our largest generator goes offline we will be able to supply demand if it is within 3 standard deviations of our prediction. This ensures we can still supply demand if any one generator goes out and our predicted demand is higher than expected. More complicated uncertainty models can fit into this framework, but developing them is beyond the scope of this work.
@@ III-A Mathematical Formulation
$\sum_{i\in G_{m}}P_{\min,i}+\sum_{i\in G_{d}}u_{i}P_{\min,i}\leq{D}_{\min,72}-3\sigma_{D}.$ (2e)
@@
$\sum_{i\in\mathcal{G}_{m}}P_{\min,i}+\sum_{\begin{subarray}{c}1\leqslant i\leqslant\tilde{m}_{\max}\\ i\in\mathcal{G}_{d}\end{subarray}}P_{\min,i}\;\geq\;D_{\min,72}+3\sigma_{D},$ (4)
@@
Eq. 2 indicates that the problem has a time decoupled objective with time coupled constraints. This approach has a potential downside of selecting a different set of generating units compared to a fully coupled approach, because the future payoffs and penalties are not accounted for at a given simulation step. It has the advantage of simplifying the computation by reducing the number of required variables.
The results of this algorithm would not be directly comparable to fully coupled approach on computational time. Since we have shown our algorithm scales sub-quadratically [27], we believe it is possible to proceed with this approach and expand it in the future to include a time-coupled objective. In this work, the runtime constraints are still time-coupled and the system would be forced to converge while fully accounting for minimum runtime or the number of starts.
@@ III-B Simulation and Results
We solve the runtime constrained problem over 2304 time periods at 5 minute intervals (8 days). This covers the peak and adjacent days of 6/17/2025-6/24/2025. The latter four days of this period was the interval in which PJM experienced the highest change of load in 2025, with the night low of 78.6 GW and the day peak of 160.2 GW, with intraday changes of 55-64 GW. We start the simulation four days before the peak, so that our 72 hour forecasts have time to impact our unit commitments before we hit the peak. The first day’s results are used as a warm-up for the system (so that it is not starting with all units off) and we exclude them from our cost function. We assume that the standard deviation of the demand forecast is $1\text{GW}$ for the full PJM system (with $924$ generators) and scaled accordingly for larger systems.
@@
Figure 2: Runtime and objective function value for RRUC algorithm including runtime constraints with 462-14784 generators. With $n$ as the number of generators, the runtime scales as $\approx O(n^{3/2})$. This is believed to be ideal in the general case [20]. The objective function per generator increases less than 5% for each doubling of the system size.
We present the results of the simulation in Fig. 2 using the same software and hardware as before. Since the reference MIP solver is too slow to be used as a benchmark for RRUC, in this study we are mainly interested in analyzing the scaling and performance of RRUC. The runtime and the objective function increase slightly faster than linearly with the number of generators. Specifically, the runtime increases by less than $3\times$ for every doubling of the system size. The $\approx O(n^{3/2})$ scaling matches best case limits, due to a bottleneck from the linear solver [20]. For the objective function, we rely on the per unit measure, or the average total cost. The per-unit objective function increases less than 5% for every doubling of the system. An ideal solver’s objective per unit would not increase with the system size. This indicates that larger systems do not result in significant diseconomies of scale and the algorithm remains cost efficient.
@@ IV-A Mathematical Formulation
As the next step, we extend the problem to not only larger and more varied thermal power systems, but also systems that have non-thermal units with intermittent characteristics. To operate such systems, we must account for system behavior at a higher resolution, which requires a sub hour granularity of unit commitment. As discussed in Section I, high temporal resolution involves another important aspect of generator behavior, ramp.
@@
Figure 3: A typical production cycle for generator $i$. When the generator is on, it can change its production power between $P_{\min,i}$ and $P_{\max,i}$, provided it does not exceed its ramping power. It can also turn off if it is close enough to $P_{\min,i}$ compared to its ability to ramp down. The only other decision can be made when the generator is off, to prepare for ramping up. During all other periods, the generator is uncontrollable.
To address the ramp constraints in the model, we use a different set of variables. Instead of the original $u_{i}$ that indicated that a generator is on or off, we allow the generator to be off, ramping up, on, or ramping down. For the model purposes, we assume that a generator can only change its status in one direction. If a generator is ramping up, it is assigned a variable $v_{i}=1$. After the required ramp time its status will switch to on. If a generator is ramping down, it is assigned a variable $w_{i}=1$. It will eventually switch off and enter the set of generators that can start. Fig. 3 shows a typical production cycle of a generator and the allowed transitions. Similarly to Section III, we use preprocessing to account for the accumulated runtime and ramping periods. The information about the duration of on, off, or ramping is stored for each generator and used as an optimization parameter.
@@
  • $\beta$ parameter we set to $0.1$. The corresponding term in the objective pushes us to choose more efficient generators for meeting the contingencies.
@@ IV-B Simulation and Results
Figure 4: Runtime and objective function value for RRUC algorithm with runtime constraints and ramping constraints for 462-14784 generators. The runtime scales as $\approx O(n^{4/3})$, where $n$ is the number of generators. The objective function per generator increases less than 2% for each doubling of the system size.
@@
The runtime and objective results are summarized in Fig. 4 and are comparable to those found in Section III using the same software and hardware. The runtime scaled roughly as $n^{4/3}$, where $n$ is the number of generators. This shows that the resulting linear systems are less complicated than those arising from a typical optimization problem ($n^{3/2}$) and that our optimization solver is close to ideal. It is well below the exponential cost of MIP solvers. The average cost of the system increased by at most 2% for each doubling of the system size. So adding extra generators does not result in significant scale cost effects.
V Unit Commitment With Non-constant Ramps
@@
Finally, we tested the suggested RRUC methodology on systems with non-constant ramp rates. The use of constant ramp rates is common in research, but they represent a simplified view of generator constraints. The ramp rates are not identical for ramping up and ramping down[3]. Ramping rates are also not constant during the ramping process, later stages of ramping are different from earlier stages[6], [31]. To account for the different ramp rates, we introduce a model extension that applies different ramp rates during different stages of generator operation.
The results were almost indistinguishable from Section IV. This is likely due to the fact that our solver is highly averse to turning units off or on, because it pays a price for it, and the generation payoff only happens in later time periods. It will only switch units on or off if it is required to meet the constraints, which happens very rarely. This motivates future work where we define our optimization objective over all time periods, instead of greedily.
@@ VI Conclusion and Further Work
We proposed an algorithm to solve unit commitment with mechanical and ramping constraints by relaxing the integer constraints and then rounding solutions so that constraints are satisfied. Our algorithm shows $\approx O(n^{4/3})-O(n^{3/2})$ time scaling in the number of generators. It retains accuracy nearly perfectly, as the problem size increases.
For a system with runtime constraints, a $2\times$ increase in the number of units results in less than $3\times$ increase in the computational time. Further, the average value of the objective function increased by less than $1.05\times$ for each $2\times$ increase in the system size, indicating that there are no diseconomies of scale associated with the increase in system size using the suggested algorithm.
When ramping constraints were added to the problem, runtime scaling was even closer to linear. The objective (cost) function per generator increased by less than 2% for each $2\times$ increase in the system size.
A switch from two ramping rates, one for ramping up and one for ramping down, to four rates for early ramping up, later ramping up, early ramping down and late ramping down, had almost no effect on the results. This may indicate that the algorithm is robust to changes in the ramp rate inputs, provided the allowed rates are fast enough.

### v2

# Fast Relax-and-Round Unit Commitment with Sub-hourly Mechanical and Ramp Constraints

@@ I Introduction
Studies directed at improving computational efficiency use comparably small simulations and one hour resolution. For example, [16] uses a bundle method on a 130-generator system, [37] uses temporal bundles and 289 generators, and [30] demonstrates security constrained unit commitment on 69 generators. There are several reasons for the use of small system sizes in unit commitment problems. First, system scaling happens along both spatial and temporal granularity. Running 1000 unit commitment scenarios for a 150-unit system remains substantially faster than running one scenario for a 1000-unit system. Similarly, running 72 time decoupled unit commitment scenarios is faster than running one time-coupled scenario for a three days outlook. Second, existing unit commitment algorithms rely on mixed integer programs (MIPs) [2]. In the worst case, solving MIPs is NP-hard and requires time exponential in the problem size, to enumerate all possibilities. For more information about the mathematical reasons for the poor scaling behavior of MIPs, see optimization-specific reviews such as [17, 20, 4, 28].
@@
Our approach only accounts for the actual costs of generation in the current period, but is biased towards turning more efficient units on. It includes intertemporal runtime constraints and, later, ramping constraints. From that perspective, the algorithm is partially temporally decoupled and cannot be directly compared to fully coupled UC algorithms. However, it still presents an advancement over earlier methods because (i) it scales sub-quadratically rather than exponentially and (ii) does not linearize the problem, making solutions more accurate. Each extended formulation is evaluated from two viewpoints: algorithmic runtime and the objective function. For the relaxed unit commitment formulation, we further analyze the stochastic characteristics of the objective function.
The study is organized as follows. Section II summarizes the original model. Section III introduces intraday unit constraints. Section IV extends the model to include ramping constraints. Section V introduces a different ramping formulation and compares its results to Section IV. Section VI summarizes the results and suggests further research directions.
@@ II-A Mathematical Formulation
  • $\sigma_{D}$: standard deviation (STD) of demand forecast
@@
The constraint Eq. 1c ensures that we supply the projected demand. The constraint Eq. 1d ensures that if our largest generator goes offline we will be able to supply demand if it is within $3\sigma_{D}$ of our prediction. This ensures we can still supply demand if any one generator goes out and our predicted demand is higher than expected. More complicated uncertainty models can fit into this framework, but developing them is beyond the scope of this work.
@@ III-A Mathematical Formulation
$\sum_{i\in G_{m}}P_{\min,i}+\sum_{i\in G_{d}}u_{i}P_{\min,i}\leq{D}_{\min,72}-\sigma_{D}.$ (2e)
Note that at the lower end, we plan for a smaller contingency ($1\sigma_{D}$) than on the higher end ($3\sigma_{D}$). This is meant to reflect the asymmetric nature of load forecasts. It is far more likely that load will be much higher than predicted (corresponding to some large unexpected load), compared to much lower (corresponding to base load being removed). This also helps provide additional buffers in the event that more than 1 generator goes out, or uncontrollable generation is much lower than expected.
@@
$\sum_{i\in\mathcal{G}_{m}}P_{\min,i}+\sum_{\begin{subarray}{c}1\leqslant i\leqslant\tilde{m}_{\max}\\ i\in\mathcal{G}_{d}\end{subarray}}P_{\min,i}\;\geq\;D_{\min,72}-\sigma_{D},$ (4)
@@
Eq. 2 indicates that the problem has a weakly time-coupled objective with strongly time-coupled constraints. This approach has a potential downside of selecting a different set of generating units compared to a fully coupled approach, because the future payoffs and penalties are not accounted for at a given simulation step. It has the advantage of simplifying the computation by reducing the number of required variables.
The results of this algorithm would not be directly comparable to fully coupled approach on computational time. Since we have shown our algorithm scales sub-quadratically [27], we believe it is possible to proceed with this approach and expand it in the future to include a fully time-coupled objective. In this work, the runtime constraints are still time-coupled and the system would be forced to converge while fully accounting for minimum runtime or the number of starts.
@@ III-B Simulation and Results
We solve the runtime constrained problem over 2304 time periods at 5 minute intervals (8 days). This covers the peak and adjacent days of 6/17/2025-6/24/2025. The latter four days of this period was the interval in which PJM experienced the highest change of load in 2025, with the night low of 78.6 GW and the day peak of 160.2 GW, with intraday changes of 55-64 GW. We start the simulation four days before the peak, so that our 72 hour forecasts have time to impact our unit commitments before we hit the peak. The first day’s results are used as a warm-up for the system (so that it is not starting with all units off) and we exclude them from our cost function. We assume that the standard deviation of the demand forecast $\sigma_{D}=1\text{GW}$ for the full PJM system (with $924$ generators) and scaled accordingly for larger systems.
@@
Figure 2: Runtime and objective function value for RRUC algorithm including runtime constraints with 462-14784 generators. With $n$ as the number of generators, the runtime scales as $\approx O(n^{1.37})$. The objective function per generator increases less than 5% for each doubling of the system size.
We present the results of the simulation in Fig. 2 using the same software and hardware as before. Since the reference MIP solver is too slow to be used as a benchmark for RRUC, in this study we are mainly interested in analyzing the scaling and performance of RRUC. The runtime and the objective function increase slightly faster than linearly with the number of generators. Specifically, the runtime increases on average about $2.6\times$ for every doubling of the system size. The $\approx O(n^{1.37})$ shows that the resulting systems are less complicated than those arising from a typical optimization problem which have $O(n^{1.5})$ scaling. For a general optimization problem, $O(n^{1.5})$ scaling is believed to be ideal, due to the bottleneck from the linear solver [20].
For the objective function, we rely on the per unit measure, or the average total cost, because our demand is scaled accordingly. The per-unit objective function increases less than 5% for every doubling of the system. It is possible we could avoid the small loss of solution quality by increasing the run time. An ideal solver’s objective per unit would not increase with the system size. However, this model is just a stepping stone to the more realistic one in  Section IV, so we will not spend time optimizing it. This indicates that larger systems do not result in significant diseconomies of scale and the algorithm remains cost efficient.
@@ IV-A Mathematical Formulation
We now extend the problem to larger and more varied thermal power systems, which have have non-thermal units with intermittent characteristics. To operate such systems, we must account for system behavior at a higher resolution, which requires a sub hour granularity of unit commitment. As discussed in Section I, high temporal resolution involves another important aspect of generator behavior, ramp.
@@
Figure 3: A typical production cycle for generator $i$ in Section IV. When the generator is on, it can change its production power between $P_{\min,i}$ and $P_{\max,i}$, provided it does not exceed its ramping power. It can also turn off if it is close enough to $P_{\min,i}$ compared to its ability to ramp down. The only other decision can be made when the generator is off, to prepare for ramping up. During all other periods, the generator is uncontrollable.
To address the ramp constraints in the model, we use a different set of variables. Instead of the original $u_{i}$ that indicated that a generator is on or off, we allow the generator to be off, preparing (for ramp up), ramping up, on, or ramping down. For the model purposes, we assume that a generator can only change its status in one direction. If a generator is ramping up or preparing, it is assigned a variable $v_{i}=1$. After the required ramp time in each stage, its status will switch to on. If a generator is ramping down, it is assigned a variable $w_{i}=1$. It will eventually switch off and enter the set of generators that can start.
Fig. 3 shows a typical production cycle of a generator and the allowed transitions. Each generator can only be in one state in a given time period, so we round ramping times from the data to whole periods. We round up to ensure that our solution is for a problem more restrictive than the actual problem, which means it is necessarily feasible. This means a generator must occupy each state for at least one time period. Similarly to Section III, we use preprocessing to account for the accumulated runtime and ramping periods. The information about the duration of on, off, or ramping is stored for each generator and used as an optimization parameter.
@@
  • $\beta$ parameter we set to $0.001$. The corresponding term in the objective pushes us to choose more efficient generators for meeting the contingencies.
@@ IV-B Simulation and Results
Figure 4: Runtime and objective function value for RRUC algorithm with runtime constraints and ramping constraints for 462-14784 generators with the Fig. 3 profile. The objective function per generator fluctuates, but does not show an increasing or decreasing trend based on the system size. The runtime scales as $\approx O(n^{1.6})$, where $n$ is the number of generators.
@@
Fig. 4 summarizes the runtime and objective results. These results are comparable to those found in Section III using the same software and hardware. The runtime scaled roughly as $n^{1.6}$, where $n$ is the number of generators. Ideal scaling is believed to be bounded below by $n^{1.5}$, due to the scaling of the underlying linear system solvers [20]. This suggests that our optimization solver has close to ideal scaling. It is well below the exponential cost of MIP solvers. The average cost of the system did not increase with the system size. So adding extra generators does not result in scale cost effects.
V Unit Commitment With Smooth Ramp Up
@@
Finally, we tested the suggested RRUC methodology on systems with smooth ramp up. The use of constant ramp rates is common in research, but they represent a simplified view of generator constraints. The ramp rates are not identical for ramping up and ramping down[3]. Ramping rates are also not constant during the ramping process, later stages of ramping are different from earlier stages[6], [31]. Further, there is no clear distinction in practice between the prepare and ramp up stages in Fig. 3. They are both part of a deterministic shift from the generator being off to on.
Figure 5: A typical production cycle for generator $i$ in Section V. When the generator is on, it can change its production power between $P_{\min,i}$ and $P_{\max,i}$, provided it does not exceed its ramping power. It can also turn off if it is close enough to $P_{\min,i}$ compared to its ability to ramp down. The only other decision can be made when the generator is off, start ramping up. During all other periods, the generator is uncontrollable.
We introduce a model extension in Fig. 5 that accounts for these discrepancies to create a more realistic model of generator states. We combine the prepare and ramp up stages into one stage where production increases quadratically from 0 to $P_{\min,i}$. This produces a more realistic, smoothed version of Fig. 3. Due to rounding up to whole time periods, this has the added effect of potentially requiring one less period for ramping up, which is closer to the actual ramp up time.
Figure 6: Runtime and objective function value for RRUC algorithm with runtime constraints and ramping constraints for 462-14784 generators, with the Fig. 5 profile. The objective function per generator fluctuates, but does not show an increasing or decreasing trend based on the system size. The runtime scales as $\approx O(n^{1.6})$.
Fig. 6 shows results using the Fig. 5 generator profile model. The results were similar to those of Section IV. There is no upward or downward trend in objective value per generator as the number of generators grows. Runtime continued to scale sub-quadratically, slightly slower than the optimal $O(n^{1.5})$.
Figure 7: Number of generators in a given state for the 462 generator simulation. Ramp up includes the preparation stage for Section IV results. Note: Only every 66th time step is shown for readability. (Top) Total number of generators in the Section IV simulation in on, ramp up, and ramp down states. Where there are 0 generators in a given state, it is omitted on the log-scale. At almost all time steps, the majority of active generators are on and controllable. (Bottom) Difference in the number of generators in a given state between the simulation in Section IV (Fig. 3 profile) and Section V (Fig. 5 profile). At almost all time steps, there is no difference.
The top part of Fig. 7 shows that most time steps have no generators ramping up or down. In $1881$ out of $2016$ time steps ($93.3\%$), all generators were on or off for the Fig. 3 generator model. In $1884$ out of $2016$ time steps ($93.5\%$), all generators were on or off for the Fig. 5 generator model. Ramp up and ramp down bursts tend to be short and include many generators. The bottom part of Fig. 7 shows the discrepancies between the solutions that both generator profiles produce. The differences in the generator states of the solutions remain large for the first $\approx 800$ time steps (equivalent to almost $3$ days), but level off. Small differences in the input can lead to very different but similar quality (similar cost) solutions, when aggregated over all time periods. This shows the non-convexity of the problem and the strength of the optimization solver, which is able to find different minima.
@@ VI Conclusion and Further Work
We proposed an algorithm to solve unit commitment with mechanical and ramping constraints by relaxing the integer constraints and then rounding solutions to satisfy the constraints. Our algorithm shows between $\approx O(n^{1.6})$ time scaling in the number of generators. As the problem size increases, it retains accuracy nearly perfectly.
For a system with only runtime constraints, a $2\times$ increase in the number of units results in about a $2.6\times$ increase in the computational time on average ($\approx O(n^{1.37})$ scaling). Further, the average value of the objective function increased by less than $1.05\times$ for each $2\times$ increase in the system size, indicating that there are some small diseconomies of scale associated with the increase in system size using the suggested algorithm.
When ramping constraints were added to the problem, runtime scaling was close to ideal. For every $2\times$ increase in problem size, runtime increased about $3\times$ on average. The objective (cost) function per generator did not increase with the problem size. It is possible that being more restrictive is beneficial for a ramp constrained formulation for objective scaling. The demand at each time step is close to the one preceding it, so a more restricted solver with ramping constraints can restrict the search on better solutions. This means that even as the problem grows, at the cost of more runtime, the solver can pinpoint similar quality solutions.
A switch to a smoother generator profile did not change the average results drastically. Objectives and their perfect scaling were similar. runtimes and their scaling were slightly better. For every $2\times$ increase in the problem size, the runtime increased less than $3\times$ on average. However, depending on the generator profile, certain time periods have substantially different units in different states. This indicates that even slightly different ramping profiles can lead to very different unit commitments, even if their objectives are similar. This highlights the non-convexity of the problem and the importance of accurate knowledge about generator profiles.